# A Kolmogorov–Arnold Neural Model for Cascading Extremes

### v1

@@ 2.3 Model Checking and Diagnostics
For evaluating the quality of the fit from the KANE model, we resort to randomized quantile residuals introduced by Dunn and Smyth, (1996). Before we present the version of these residuals for our model, we recall that their basic principle is as follows: For any nonparametric or parametric model, based on a continuous distribution function, $F(x)=P(X\leq x)$, it follows that $R\equiv\Phi^{-1}(F(X))\sim N(0,1)$, when $X\sim F$, where $\Phi^{-1}$ is the standard normal quantile function. Hence, provided that $F$ is a sensible model for the data $X$, then the residuals $R$ are a normal distributed. As noted by Dunn and Smyth,, the same principle can be easily extended to a discrete model $F$ via jittering. Since for each fixed $u$, our POC surface formulation can be understood as binary response regression model based on $D_{n}$, it follows that its Dunn–Smyth residuals are defined as
@@
Loosely, the $R_{i}$ are a jittered version of the ‘binary response’ (i.e., follow-up event $\delta_{i}$), with the level of jittering being controlled by the ‘probability of success’ (i.e., probability of cascade $\hat{\alpha}(\mathbf{x}_{i})$). We recommend reporting at least 10 trajectories of the randomized quantile residuals, noting that Dunn and Smyth, (1996) suggested that “four realizations of the quantile residual” would already be sufficient. These realizations will be presented in a QQ-boxplot along with the confidence bands stemming from the normal reference model (that keeps in mind that the $R_{i}$ are normal distributed provided that the fit is sensible); the QQ-boxplot was recently introduced by Rodu and Kafadar, (2021), and consists of a graphical tool that merges into a single chart a boxplot and a QQ-plot.
@@ 3.1 Multi-Trigger Systems
In practice, multiple competing incidents can result in the follow-up event. To address this, we define a multi-trigger system where we consider $\mathcal{Y}_{1},\dots,\mathcal{Y}_{K}$ as a sequence of identically distributed random fields with
$\mathcal{Y}_{k}=\{Y_{k,\mathbf{x}}:\mathbf{x}\in[0,1]^{d}\},\qquad k=1,\dots,K.$
@@
$\alpha(\mathbf{x})=\lim_{u\to y^{*}}P(I_{u,\mathbf{x}}=1\mid\,Y_{1,\mathbf{x}}>u\vee\cdots\vee Y_{K,\mathbf{x}}>u).$ (10)
This formula simplifies to the single-trigger case in (6) by defining $Y_{\mathbf{x}}=\min\{Y_{1,\mathbf{x}},\dots,Y_{K,\mathbf{x}}\}$, and hence the theory and methods from Section 2 readily apply to this context as well. The identically distributed assumption for $\mathcal{Y}_{1},\dots,\mathcal{Y}_{K}$ is for convenience, as in practice random processes can be normalized to a common scale (say, unit Fréchet margins) using the probability integral transform.
@@ 3.2 Categorical, Ordinal, and Continuous Follow-up Events
is a random field with multinoulli marginal distribution. Here, $\mathbf{I}_{u,\mathbf{x}}=(I_{u,\mathbf{x}}^{(1)},\dots,I_{u,\mathbf{x}}^{(J)})^{\top}\in\{0,1\}^{J}$ and $J$ is the number of categories; see Murphy, (2012, Section 2.3.2) for basics on the multinoulli distribution. The proposed cascading probability extends to this context as follows.
@@
Let $\mathbf{x}=(x_{1},\dots,x_{d})^{\top}\in[0,1]^{d}$ and $J\in\mathbb{N}$. The $j$th-category probability of cascade surface is defined as
@@
for $j\in\{1,\dots,J\}$, where $\mathbf{I}=\{\mathbf{I}_{u,\mathbf{x}}:(u,\mathbf{x})\in\mathbb{R}\times[0,1]^{d}\}$ is a multinoulli process and $\{Y_{\mathbf{x}}:\mathbf{x}\in[0,1]^{d}\}$ is a random field.
In terms of fitting the model, two adjustment are needed for the architecture discussed in Section 2. Firstly, we need an activation function for the $g$-layer that maps the Euclidean space to the unit simplex; that is, $\textbf{g}=(g^{(1)},\dots,g^{(J)})$, where $g^{(j)}:\mathbb{R}^{d}\to[0,1]$, such that $\sum_{j=1}^{J}g^{(j)}(\mathbf{x})=1$. A natural candidate for this task is the softmax activation function, in which case the $g$-layer becomes,
$g^{(j)}(\mathbf{x})=\frac{\exp(x_{j})}{\sum_{j=1}^{J}\exp(x_{j})}.$ (12)
@@
where $(\delta_{i,1},\dots,\delta_{i,J})^{\top}$ is a realization of $\mathbf{I}_{u,\mathbf{x}_{i}}$.
@@
The Frank–Hall approach begins by transforming a $J$-class ordinal problem into $J-1$ binary classification problems. Specifically, an ordinal attribute $A$ with ordered values $C_{1},\dots,C_{J}$ is converted into $J-1$ binary attributes of the form $A>C_{j}$, for $j=1,\dots,J-1$; for example, on the above-mentioned $J=5$ hurricane problem, this would imply converting the original ordinal problem into four binary attributes based on
@@ 4.1 Artificial Data Generating Processes and Preliminary Experiments
[…]
@@
$\alpha^{(1)}_{\mathbf{I}}(\mathbf{x})&=\lim_{u\to\infty}\frac{m_{\text{C1}}(\mathbf{x};u)}{M_{u}}=\frac{\Phi(\mathbf{x};\boldsymbol{0},\boldsymbol{I}_{2})}{M},\\ \alpha^{(2)}_{\mathbf{I}}(\mathbf{x})&=\lim_{u\to\infty}\frac{m_{\text{C2}}(\mathbf{x};u)}{M_{u}}=\frac{0.4\exp(-x_{1})\cos\{2\pi x_{2}\}+0.5}{M},\\ \alpha^{(3)}_{\mathbf{I}}(\mathbf{x})&=\lim_{u\to\infty}\frac{m_{\text{C3}}(\mathbf{x};u)}{M_{u}}=\frac{0.8x_{2}\sin^{2}(\pi x_{1})}{M},\\ \end{aligned}\end{cases}$
@@
The results from a single-sample experiment are shown in Figure 2. As can be seen from the latter figures, the proposed methods satisfactorily recover the true POC in all scenarios; these findings should be considered provisional, as they come from a single-sample experiment, with the Monte Carlo performance to be evaluated next in Section 4.2.
Some final comments on computing and implementations are in order. The neural model was fitted using the LBFGS optimizer (Pytlak,, 2008, Chapter 5). For binary classification scenarios, the sigmoid activation function (9) was specified for the $g$ layer to ensure that the output values remain within the unit interval, representing valid probabilities. In cases of multi-category classification, the softmax activation function in (12) was used. The optimization process was carried out for 100 steps for each scenario, with $m=2$ and $p=3$, as cubic splines are continuous and have continuous first and second derivatives at each knot.
@@
4.2 Monte Carlo Evidence
[…]
Table 1: MISE values across scenarios and sample sizes.
[…]
5 00050005\,0005 000 10 0001000010\,00010 000 15 0001500015\,00015 000
[…]
A2 $1.501\times 10^{-4}$ $1.541\times 10^{-4}$ $1.478\times 10^{-4}$
B1 $1.021\times 10^{-4}$ $8.014\times 10^{-5}$ $7.722\times 10^{-5}$
B2 $2.669\times 10^{-3}$ $2.626\times 10^{-3}$ $2.625\times 10^{-3}$
C ($j=1$) $9.465\times 10^{-4}$ $9.055\times 10^{-4}$ $8.958\times 10^{-4}$
C ($j=2$) $1.166\times 10^{-3}$ $1.130\times 10^{-3}$ $1.106\times 10^{-3}$
C ($j=3$) $1.725\times 10^{-4}$ $1.185\times 10^{-4}$ $9.428\times 10^{-5}$
[…]
Finally, in Table 4.2 we report the Monte Carlo MISE (Mean Integrated Squared Error), where
Here, $\hat{\alpha}_{I}(\mathbf{x})$ is the KANE estimate of POC and $[0,1]^{d}$ is the space of the features (i.e., $d=1$ for Scenarios A1–A2, and $d=2$ for Scenarios B1–B2 and C). As can be seen from Table 4.2, the proposed approach has appealing frequentist properties from a numerical viewpoint, in the sense that as the sample size increases, MISE decreases.
In the supplementary materials, we document a battery of additional Monte Carlo experiments, including analyses of dependent features, an evaluation of the deep model, and a comparison of the Frank–Hall ordinal model with the softmax approach. Among other findings, these additional results further validate the approach’s performance, show that performance is better in data-rich regions of the covariate space, and that the canonical three-layer model may perform as well as, if not better than, its deep counterpart.
@@ 5.1 Earthquake–Tsunami Data
The trigger events are thresholded at $u=F_{Y_{\mathbf{x}}}^{-1}(0.95)$, resulting in 167 observations of earthquakes with magnitudes greater than 7.9 on the Richter scale, along with the corresponding indicators of tsunami occurrence. The fitted POC surfaces were modeled according to Section 2.2, with a three-layer KANE framework following the setup from Section 4, and considering the features $\mathbf{x}=(\text{latitude, longitude, depth})^{\top}$.
As can be seen from Fig. 5, the fitted surface tends to be particularly pronounced, for example in the coast of Chile, California, and the Pacific ring in comparison with the Caribbean; the same figure also illustrates that POC tends to be low at greater depths (e.g., 85th percentile), hence suggesting that earthquakes which are deeper would less likely lead to a cascading tsunami event. All in all, the usual extrapolation disclaimer applies. Interpretations should account for the fact that the quality of the fits is more reliable in regions with greater data availability, while their accuracy should be questioned in regions with limited data. Finally, the QQ-boxplot of the randomized residuals, shown in Fig. 7, indicates an overall good fit; further trajectories in the supplementary materials indicate a similar finding.
@@ 5.2 Tropical Cyclone–Sea Surface Temperature Data
[Uncaptioned image] [Uncaptioned image]
@@
Keeping in mind the ordinal nature of the data, the fitted POC surfaces were modeled according to the Frank–Hall variant of our model (Section 3), with a three-layer KANE framework, and considering the features $\mathbf{x}=(\text{latitude, longitude})^{\top}$. Among other insights, the fitted surfaces suggest that high sea surface temperatures tend be associated with more extreme tropical storms in parts of Gulf of Mexico. Aside from this, the analysis also reveals that the probability of cascade remains relatively small for tropical depression and hurricanes. Finally, the QQ-boxplot of the randomized residuals is shown in Fig. 7 and suggests an overall good fit of the model; further trajectories, presented in the supplementary materials, provide comparable evidence.
@@
Figure 7: QQ-boxplots of Dunn–Smyth residuals for empirical illustrations.
@@ 6 Implications and Future Directions
This paper develops a novel statistical framework to address the growing concern of cascading extreme events—such as earthquakes triggering tsunamis or heatwaves sparking wildfires, which in turn lead to further losses. By integrating EVT and AI, the proposed methods aim to enable extrapolation beyond observed data to capture and model complex chains of extreme events. Although straightforward, the additional ‘Natural Enforcement’ of KAN discussed in this work may be valuable for other applications. For instance, it could be relevant when modeling covariate-adjusted versions of other extreme value parameters, such as $\bar{\chi}$ (Coles,, 2001, Chapter 8), which is constrained to the range $(-1,1]$. This constraint could similarly be enforced naturally within our $g$-layer.
While KAN models have been claimed to be explainable (Liu et al.,, 2024), one might consider the possibility of fitting POC surfaces using a more conventional explainable statistical model. If explainability is a priority, the POC surfaces introduce above can be formulated as an additive generalized linear models along the same lines as in Lee et al., (2024).
We close the paper with some final comments on future research. In practice, either the follow-up or the trigger event may themselves functional in the sense of FDA (Functional Data Analysis) (Horváth and Kokoszka,, 2012; Kokoszka,, 2017). For example, we may observe $I_{u,\mathbf{x}}=1$ for $\mathbf{x}$ over a continuum rather than over a point $\mathbf{x}\in[0,1]^{d}$. As a concrete instance of this, in the earthquake data application, a follow-up event could represent the full region $S\subset[0,1]^{2}$ affected by the tsunami, in which case $I_{u,\mathbf{x}}=1$ for all $\mathbf{x}\in S$. While our theory also accommodates this framework, further investigation is needed to incorporate such functional events into the inferences in a fully FDA-aligned fashion. Secondly, while there are fundamental differences between our approach and Hawkes processes, there may be interest in combining our framework with the cross-excitation of multiple point processes introduced by Hawkes, (1971).
Finally, while the proposed approach targets two-chain setups involving a trigger and follow-up extreme event, it may set the stage for developments that accommodate extremal chains of random length ($k$), and also factoring in the role of the chain’s pathways. TDMs (Tail Dependence Matrices) (Embrechts et al.,, 2016) may be used to govern the pathway of the extremal cascade; in this setting the order of chained events could be dictated by a random permutation of $\{1,\dots,k\}$, based on transition probabilities for the TDMs between extremal events. This would allow for modelling and learning from the data the length or size of the extremal cascade ($k$) and the distribution of the pathway taken by it (say, $4\to 1\to 2\to 3$ or $3\to 2\to 1\to 4$ as two examples of realizations of the chain of extremes).

### v2

@@ 2.3 Model Checking and Diagnostics
For evaluating the quality of the fit from the KANE model, we resort to randomized quantile residuals introduced by Dunn and Smyth, (1996). Before we present the version of these residuals for our model, we recall that their basic principle is as follows: For any nonparametric or parametric model, based on a continuous distribution function, $F(x)=P(X\leq x)$, it follows that $R\equiv\Phi^{-1}(F(X))\sim N(0,1)$, when $X\sim F$, where $\Phi^{-1}$ is the standard normal quantile function. Hence, provided that $F$ is a sensible model for the data $X$, then the residuals $R$ are normally distributed. As noted by Dunn and Smyth,, the same principle can be easily extended to a discrete model $F$ via jittering. Since for each fixed $u$, our POC surface formulation can be understood as binary response regression model based on $D_{n}$, it follows that its Dunn–Smyth residuals are defined as
@@
Informally, the $R_{i}$ are a jittered version of the ‘binary response’ (i.e., follow-up event $\delta_{i}$), with the level of jittering being controlled by the ‘probability of success’ (i.e., probability of cascade $\hat{\alpha}(\mathbf{x}_{i})$). We recommend reporting at least 10 trajectories of the randomized quantile residuals, noting that Dunn and Smyth, (1996) suggested that “four realizations of the quantile residual” would already be sufficient. These realizations will be presented in a QQ-boxplot along with the confidence bands stemming from the normal reference model (that keeps in mind that the $R_{i}$ are normally distributed provided that the fit is sensible); the QQ-boxplot was recently introduced by Rodu and Kafadar, (2021), and consists of a graphical tool that merges into a single chart a boxplot and a QQ-plot.
@@ 3.1 Multi-Trigger Systems
In practice, multiple competing incidents can result in the follow-up event. To address this, we define a multi-trigger system where we consider $\mathcal{Y}_{1},\dots,\mathcal{Y}_{\mathcal{K}-1}$ as a sequence of identically distributed random fields with
$\mathcal{Y}_{k}=\{Y_{k,\mathbf{x}}:\mathbf{x}\in[0,1]^{d}\},\qquad k=1,\dots,\mathcal{K}.$
@@
$\alpha(\mathbf{x})=\lim_{u\to y^{*}}P(I_{u,\mathbf{x}}=1\mid\,Y_{1,\mathbf{x}}>u\vee\cdots\vee Y_{\mathcal{K},\mathbf{x}}>u).$ (11)
This formula simplifies to the single-trigger case in (6) by defining $Y_{\mathbf{x}}=\min\{Y_{1,\mathbf{x}},\dots,Y_{\mathcal{K},\mathbf{x}}\}$, and hence the theory and methods from Section 2 readily apply to this context as well. The identically distributed assumption for $\mathcal{Y}_{1},\dots,\mathcal{Y}_{\mathcal{K}}$ is for convenience, as in practice random processes can be normalized to a common scale (say, unit Fréchet margins) using the probability integral transform.
@@ 3.2 Categorical, Ordinal, and Continuous Follow-up Events
is a random field with multinoulli marginal distribution. Here, $\mathbf{I}_{u,\mathbf{x}}=(I_{u,\mathbf{x}}^{(1)},\dots,I_{u,\mathbf{x}}^{(J)})^{T}\in\{0,1\}^{J}$ and $J$ is the number of categories; see Murphy, (2012, Section 2.3.2) for basics on the multinoulli distribution. The proposed cascading probability extends to this context as follows.
@@
Let $\mathbf{x}=(x_{1},\dots,x_{d})^{T}\in[0,1]^{d}$ and $J\in\mathbb{N}$. The $j$th-category probability of cascade surface is defined as
@@
for $j\in\{1,\dots,J\}$, given that the limit exists, where $\mathbf{I}=\{\mathbf{I}_{u,\mathbf{x}}:(u,\mathbf{x})\in\mathbb{R}\times[0,1]^{d}\}$ is a multinoulli process and $\{Y_{\mathbf{x}}:\mathbf{x}\in[0,1]^{d}\}$ is a random field.
In terms of fitting the model, two adjustments are needed for the architecture discussed in Section 2. Firstly, we need an activation function for the $g$-layer that maps the Euclidean space to the unit simplex; that is, $\textbf{g}=(g^{(1)},\dots,g^{(J)})$, where $g^{(j)}:\mathbb{R}^{d}\to[0,1]$, such that $\sum_{j=1}^{J}g^{(j)}(\mathbf{x})=1$. A natural candidate for this task is the softmax activation function, in which case the $g$-layer becomes,
$g^{(j)}(\mathbf{x})=\frac{\exp(x_{j})}{\sum_{i=1}^{J}\exp({x_{i}})}.$ (13)
@@
where $(\delta_{i,1},\dots,\delta_{i,J})^{T}$ is a realization of $\mathbf{I}_{u,\mathbf{x}_{i}}$.
@@
The Frank–Hall approach begins by transforming a $J$-class ordinal problem into J1J-1 binary classification problems. Specifically, an ordinal attribute $A$ with ordered values $C_{1},\dots,C_{J}$ is converted into J1J-1 binary attributes of the form $A>C_{j}$, for $j=1,\dots,J-1$; for example, on the above-mentioned $J=5$ hurricane problem, this would imply converting the original ordinal problem into four binary attributes based on
@@ 4.1 Artificial Data Generating Processes and Preliminary Experiments
This section presents the simulation scenarios used to evaluate the proposed methods, along with a series of single-sample experiments. We assess performance in three settings by simulating n=10 000n=10\,000 observations per scenario, with artificial data generated from the Bernoulli process, follow-up events, and features as described below. In each case, we threshold the response at its 95% quantile ($u$), hence retaining only those values where $Y_{x}>u$.
@@
[…]
@@
The results from a single-sample experiment are shown in Fig. 2. As can be seen from the latter figures, the proposed methods satisfactorily recover the true POC in all scenarios; these findings are provisional, since they are based on a single-sample experiment; a Monte Carlo evaluation follows in Section 4.2.
Some final comments on computing and implementations are in order. The neural model was fitted using the LBFGS optimizer (Pytlak,, 2008, Chapter 5). For binary classification scenarios, the sigmoid activation function (10) was specified for the $g$-layer to ensure that the output values remain within the unit interval, representing valid probabilities. In cases of multi-category classification, the softmax activation function in (13) was used. The optimization process was carried out for 100 steps for each scenario. A key consideration in model fitting was the selection of $p$ and $m$; for the first decision, we relied on the properties of cubic splines and chose $p=3$, as cubic splines are continuous and have continuous first and second derivatives. The number of knots was set to $m+1=3$, with two knots placed at the boundaries and one interior knot to provide flexibility; this represents the minimal configuration that yielded satisfactory performance in our experiments.
4.2 Monte Carlo Evidence
This section reports the main findings of a Monte Carlo simulation study, where we repeat $500$ times the single-sample experiment from the previous section. We consider three sample sizes n=5000,10 000,n=5000,10\,000, and 15 00015\,000. Figure 3 shows the Monte Carlo means of the simulation study, indicating strong performance across different scenarios, as anticipated by Fig. 2.
Table 1: MISE values across scenarios and sample sizes.
          Scenario       Sample Size
  5 0005\,000 10 00010\,000 15 00015\,000
      A1 $9.112\times 10^{-6}$ $3.706\times 10^{-6}$ $7.759\times 10^{-7}$
    A2 $1.501\times 10^{-4}$ $1.541\times 10^{-4}$ $1.478\times 10^{-4}$
    B1 $1.021\times 10^{-4}$ $8.014\times 10^{-5}$ $7.722\times 10^{-5}$
    B2 $69\times 10^{-3} 2.626$$26\times 10^{-3}$ $2.625\times 10^{-3}$
    C ($j=1$) $9.465\times 10^{-4}$ $9.055\times 10^{-4}$ $8.958\times 10^{-4}$
    C ($j=2$) $166\times 10^{-3} 1.1$$30\times 10^{-3} 1.1$$106\times 10^{-3}$
    C ($j=3$) $725\times 10^{-4} 1.1$$185\times 10^{-4}$ $9.428\times 10^{-5}$
Finally, in Table 1 we report the Monte Carlo MISE (Mean Integrated Squared Error), where
Here, $\hat{\alpha}_{I}(\mathbf{x})$ is the KANE estimate of POC and $0,1]^{d}$ is the space of the features (i.e., $d=1$ for Scenarios A1–A2, and $d=2$ for Scenarios B1–B2 and C). As can be seen from Table 1, the proposed approach has appealing frequentist properties from a numerical viewpoint, in the sense that as the sample size increases, MISE decreases.
In the supplementary materials, we document a battery of additional Monte Carlo experiments, including analyses of dependent features, an evaluation of the deep model, higher-dimensional features, and a comparison of the Frank–Hall ordinal model with the softmax approach. Among other findings, these additional results further validate the approach’s performance, show that performance is better in data-rich regions of the covariate space, and that the canonical three-layer model may perform as well as, if not better than, its deep counterpart, which also requires careful selection of the number of layers to avoid overfitting.
Moreover, the canonical three-layer KANE is computationally efficient and stable, consistently converging in under one second per scenario on a standard laptop (Apple M3) for $p=2$. However, as noted in the supplementary materials, increasing the feature dimension to $p>5$ leads to a substantial rise in computational time. Finally, we note that estimators for the POC surface can also be constructed using Generalized Additive Models (GAM) and Multi-Layer Perceptrons (MLP). Numerical evidence in the supplementary materials suggests that KANE and MLP yield comparable POC surface estimates, with both outperforming GAM.
@@ 5.1 Earthquake–Tsunami Data
The trigger events are thresholded at $u=F_{Y_{\mathbf{x}}}^{-1}(0.95)$, resulting in 167 observations of earthquakes with magnitudes greater than 7.9 on the Richter scale, along with the corresponding indicators of tsunami occurrence. The fitted POC surfaces were modeled according to Section 2.2, with a three-layer KANE framework following the setup from Section 4, and considering the features $\mathbf{x}=(\text{latitude, longitude, depth})^{T}$.
As can be seen from Fig. 5, the fitted surface tends to be particularly pronounced, for example in the coast of Chile, California, and the Pacific ring in comparison with the Caribbean; the same figure also illustrates that POC tends to be low at greater depths (e.g., 85th percentile), hence suggesting that earthquakes which are deeper would be less likely to lead to a cascading tsunami event. All in all, the usual extrapolation disclaimer applies. Interpretations should account for the fact that the quality of the fits is more reliable in regions with greater data availability, while their accuracy should be questioned in regions with limited data. Finally, the QQ-boxplot of the randomized residuals, shown in Fig. 7, indicates an overall good fit; further trajectories in the supplementary materials indicate a similar finding.
@@ 5.2 Tropical Cyclone–Sea Surface Temperature Data
Refer to caption Refer to caption
@@
Keeping in mind the ordinal nature of the data, the fitted POC surfaces were modeled according to the Frank–Hall variant of our model (Section 3), with a three-layer KANE framework, and considering the features $\mathbf{x}=(\text{latitude, longitude})^{T}$. Among other insights, the fitted surfaces suggest that high sea surface temperatures tend to be associated with more extreme tropical storms in parts of the Gulf of Mexico. Aside from this, the analysis also reveals that the probability of cascade remains relatively small for tropical depression and hurricanes. Finally, the QQ-boxplot of the randomized residuals is shown in Fig. 7 and suggests an overall good fit of the model; further trajectories, presented in the supplementary materials, provide comparable evidence.
@@
Figure 7: $Q$-boxplots of Dunn–Smyth residuals for empirical illustrations.
@@ 6 Implications and Future Directions
This paper develops a novel statistical framework to address the growing concern of cascading extreme events—such as earthquakes triggering tsunamis or heatwaves sparking wildfires, which in turn lead to further losses. By integrating EVT and AI, the proposed methods aim to enable extrapolation beyond observed data to capture and model complex chains of extreme events. Although straightforward, the additional ‘Natural Enforcement’ of KAN discussed in this work may be valuable for other applications. For instance, it could be relevant when modeling covariate-adjusted versions of other extreme value parameters, such as $\bar{\chi}$ (Coles,, 2001, Chapter 8), which is constrained to the range (1,1](-1,1]. This constraint could similarly be enforced naturally within our $g$-layer.
While KAN models have been claimed to be explainable (Liu et al.,, 2024), one might consider the possibility of fitting POC surfaces using a more conventional explainable statistical model. If explainability is a priority, the POC surfaces introduced above can be formulated as an additive generalized linear models along the same lines as in Lee et al., (2024).
We close the paper with some final comments on future research. In practice, either the follow-up or the trigger event may themselves be functional in the sense of FDA (Functional Data Analysis) (Horváth and Kokoszka,, 2012; Kokoszka,, 2017). For example, we may observe $I_{u,\mathbf{x}}=1$ for $\mathbf{x}$ over a continuum rather than over a point $\mathbf{x}\in[0,1]^{d}$. As a concrete instance of this, in the earthquake data application, a follow-up event could represent the full region $S\subset[0,1]^{2}$ affected by the tsunami, in which case $I_{u,\mathbf{x}}=1$ for all $\mathbf{x}\in S$. While our theory also accommodates this framework, further investigation is needed to incorporate such functional events into the inferences in a fully FDA-aligned fashion. Secondly, while there are fundamental differences between our approach and Hawkes processes, there may be interest in combining our framework with the cross-excitation of multiple point processes introduced by Hawkes, (1971).
Finally, while the proposed approach targets two-chain setups involving a trigger and follow-up extreme event, it may set the stage for developments that accommodate extremal chains of random length ($k$), and also account for the role of the chain’s pathways. TDMs (Tail Dependence Matrices) (Embrechts et al.,, 2016) may be used to govern the pathway of the extremal cascade; in this setting the order of chained events could be dictated by a random permutation of $\{1,\dots,k\}$, based on transition probabilities for the TDMs between extremal events. This would allow for modelling and learning from the data the length or size of the extremal cascade ($k$) and the distribution of the pathway taken by it (say, $4\to 1\to 2\to 3$ or $3\to 2\to 1\to 4$ as two examples of realizations of the chain of extremes).